Depreciation by Three Methods; Partial Years Perdue Company purchased equipment on April 1 for $32,670. The...

  1. Depreciation by Three Methods; Partial Years

    Perdue Company purchased equipment on April 1 for $32,670. The equipment was expected to have a useful life of three years, or 7,020 operating hours, and a residual value of $1,080. The equipment was used for 1,300 hours during Year 1, 2,500 hours in Year 2, 2,100 hours in Year 3, and 1,120 hours in Year 4.

    Required:

    Determine the amount of depreciation expense for the years ended December 31, Year 1, Year 2, Year 3, and Year 4, by (a) the straight-line method, (b) units-of-output method, and (c) the double-declining-balance method.

    Note: FOR DECLINING BALANCE ONLY, round the multiplier to four decimal places. Then round the answer for each year to the nearest whole dollar.

    a. Straight-line method

    Year Amount
    Year 1 $
    Year 2 $
    Year 3 $
    Year 4 $

    b. Units-of-output method

    Year Amount
    Year 1 $
    Year 2 $
    Year 3 $
    Year 4 $

    c. Double-declining-balance method

    Year Amount
    Year 1 $
    Year 2 $
    Year 3 $
    Year 4 $
Answer #1

Requirement a

Straight line Method
A Cost $        32,670.00
B Residual Value $           1,080.00
C=A - B Depreciable base $        31,590.00
D Life [in years left ]                             3
E=C/D Annual SLM depreciation $        10,530.00
Depreciation schedule-Straight line method
Year Book Value Depreciation expense Accumulated Depreciation Ending Book Value
1 $          32,670.00 $                 7,898 $         7,897.50 $      24,772.50
2 $          24,772.50 $              10,530 $      18,427.50 $      14,242.50
3 $          14,242.50 $              10,530 $      28,957.50 $         3,712.50
4 $            3,712.50 $                 2,633 $      31,590.00 $         1,080.00

Requirement b

Units of output method
A Cost $        32,670.00
B Residual Value $           1,080.00
C=A - B Depreciable base $        31,590.00
D Usage in units(in Hours) 7020
E Depreciation per hour $                   4.50
Depreciation schedule-Units of Output
Year Book Value Usage Depreciation expense Ending Book Value Accumulated Depreciation
1 $          32,670.00 1300 $         5,850 $      26,820.00 $         5,850.00
2 $          26,820.00 2500 $      11,250 $      15,570.00 $      17,100.00
3 $          15,570.00 2100 $         9,450 $         6,120.00 $      26,550.00
4 $            6,120.00 1120 $         5,040 $         1,080.00 $      31,590.00

Requirement c

Double declining Method
A Cost $        32,670.00
B Residual Value $           1,080.00
C=A - B Depreciable base $        31,590.00
D Life [in years] 3
E=C/D Annual SLM depreciation $        10,530.00
F=E/C SLM Rate 33.33%
G=F x 2 DDB Rate 66.6667%
Depreciation schedule-Double declining
Year Beginning Book Value Depreciation rate Depreciation expense Accumulated Depreciation Ending Book Value
1 $          32,670.00 66.6667% $            16,335 $      16,335.01 $      16,334.99
2 $          16,334.99 66.6667% $            10,890 $      27,225.01 $         5,444.99
3 $            5,444.99 66.6667% $               3,630 $      30,855.00 $         1,815.00
4 $            1,815.00 66.6667% $                  735 $      31,590.00 $         1,080.00